# Cosmic Collider Gravitational Waves sourced by Right-handed Neutrino production from Bubbles: Testing Seesaw, Leptogenesis and Dark Matter

### v1

@@ VI Viable UV-complete Particle Theory: Multi-Majoron Model
where $\rm{Br}_{\varphi_{1}\rightarrow N_{3}}$ is the branching ratio of $\varphi_{1}$ into the RHN $N_{3}$ state, and $E_{\max}$ is given in Eq. (43). Note due to the mixing between $\phi_{1}$ and $\phi_{2}$, we have to diagonalize the mass matrix coming from Eq. 76 in order to find $m_{\varphi_{1}}$. Following Ref. DiBari:2023mwu we find the corrected mass
@@
Note after $\varphi_{1}$ P.T., $N_{3}$ gets a mass $M_{3}=y_{\phi_{1}}v_{1}/2$. Hence in order for $\varphi_{1}$ to decay into a pair of $N_{3}$ one requires
@@
Hence for our benchmark points with $v_{1}=10^{14}$ GeV, we can neglect the vev correction and simply get an upper bound $y_{3}\lesssim 0.45$. This leads to a corresponding upper bound on $M_{3}$ while calculating different relic density and yields. We take CP violation to be $\epsilon_{\rm max}\sim\mathcal{O}(1)$ [^18] For instance in the resonant leptogenesis $\epsilon_{\rm CP}$ can be made very large Pilaftsis:2003gt.. Note, in our case $\lambda_{2}/\zeta\lesssim\mathcal{O}(10^{-3})$ hence we can safely assume $\varphi_{2}$ self production from $\varphi_{2}$ bubble collision remains subdominant to $\varphi_{1}$ production. By inspecting Eq. 47 we see the produced $N_{3}$ yield from $\varphi_{2}$ collision is $\propto\ln\left(\frac{2\rm E_{max}}{M_{3}}\right)$ whereas the produced $\varphi_{2}$ yield is $\propto\left[\ln\left(\frac{\rm E_{max}}{m_{\varphi_{1}}}\right)+\frac{8\pi^{2}v_{1}^{2}}{m_{\varphi_{1}}^{2}}\right]$. We note for our benchmark points, $\frac{8\pi^{2}v_{1}^{2}}{m_{\varphi_{1}}^{2}}\gg\ln\left(\frac{\rm E_{max}}{m_{\varphi_{1}}}\right),\ln\left(\frac{2\rm E_{max}}{M_{3}}\right)$, hence the produced $N_{3}$ density always remain subdominant to the produced $\varphi_{1}$ density. Finally, we point out that the mass of the produced $\varphi_{1}$ particles, $m_{\varphi_{1}}\sim\sqrt{2\lambda_{1}}v_{1}\gg v_{2}\sim T_{*}$ for our benchmark and hence the thermal contribution for $\varphi_{1}$ yield is negligible at production.
@@
The corresponding values of the DM relic and BAU for the benchmark points are summarized in Table. 3.Using Eqs. (48) and (49), we find that in the RHN dark matter scenario the observed relic abundance is reproduced for RHN masses in the range $M_{3}\sim 10^{5}$$10^{7}$ TeV. For leptogenesis, the baryon yield $\mathcal{Y}_{B}$ is computed using Eq. (66), assuming $\epsilon_{\rm max}\sim\mathcal{O}(1)$ Pilaftsis:2003gt. For $BP_{Y}$ and $BP_{Z}$ we produce the observed baryon yield given in Eq. 64 for RHN mass $10^{7}-10^{8}$ TeV. On the other hand, for $BP_{X}$, even for the maximum $M_{3}$ allowed by the bound Eq. 87, roughly half of the required asymmetry is produced. This happens due to the relatively low P.T. temperature for $BP_{X}$.
@@
The model also allows for a co-genesis scenario through the inclusion of an additional interaction term of the form $\mathcal{L}\sim y_{\chi}\bar{\chi}\phi_{2}N_{3}$ in Eq. 75. The additional dark sector fermion $\chi$ is charged under $U(1)_{N}\times U(1)_{B-L}$ such that this operator is invariant under the extended multi-Majoron symmetry. The fermion $\chi$ then serves as the DM candidate. The analysis closely parallels that presented in Sec. V. The inclusion of this extra fermionic degree of freedom does not significantly affect the effective thermal potential, since both the thermal and one-loop corrections from fermions are relatively small Salvio:2024upo. Consequently, the phase transition parameters may be taken to be identical to those computed previously.
@@
In the co-genesis scenario, the dark matter relic abundance is reproduced for $\text{Br}\chi m\chi\sim 5$$50$ GeV. For $\text{Br}_{\chi}\sim 1$, this corresponds to heavy dark matter with $m_{\chi}\sim 10$ GeV; however, in this limit $\text{Br}_{L}\ll 1$ and the resulting baryon asymmetry is negligible. Conversely, for $\text{Br}_{\chi}\sim 0.1$, the observed baryon asymmetry can be obtained simultaneously with a light dark matter candidate, $m_{\chi}\sim\mathcal{O}(100)$ GeV.
@@
Figure 15: Parameter scan for a FOPT in the multi-Majoron model in $v_{2}-M$ plane for differnt choices of $\zeta$ and $\lambda_{2}$. Shown are parameter-space points that realize a successful FOPT and produce GW signals with SNR >10>10 at ET and BBO, arising from both bubble collisions and $\varphi_{1}$ production. Regions compatible with successful dark matter production, leptogenesis, and co-genesis are indicated, together with the corresponding values of the RHN mass $M_{3}$. $\kappa_{\varphi_{1}}\sim\lambda\sim\mathcal{O}(1)$ and $\lambda_{1}=0.001$ is taken for all the plots.
@@ VII Discussion & Conclusion
  • A non-thermal population of RHNs can be generated from bubble collisions during a FOPT. If the lightest among them is stable then they may constitute the entire DM relic abundance for masses as low as $m_{DM}\gtrsim 10^{6}\ \mathrm{GeV}$ (Fig. 5). Higher transition temperatures correspond to lighter viable DM masses (see Eqs. 48 and 49). The characteristic $f^{1}$ slope from RHN production will be detectable in LISA for $m_{DM}\sim 10^{10}\ \mathrm{GeV}$, SKA for $m_{DM}\sim 10^{14}\ \mathrm{GeV}$ and ET and BBO for $m_{DM}\sim 10^{6}\ \mathrm{GeV}$ (see Fig. 6). The novel correlation between the DM relic, mass and the GW amplitude can be understood from Eqs. (34-36,47-49) via $\alpha,\ \beta/H,\ T_{*}$ and $y_{\rm DM}$.
@@ Appendix B Processes leading to Washout of Asymmetries
Here we note down a few washout processes related to our scenario discussed in sec IV. The expressions for all scattering cross sections are approximate and we have dropped the contributions of any thermal masses or subleading logarithmic pieces cataldi2024leptogenesisbubblecollisions -
@@ Appendix C RHNs and Radiation Domination
Where in deriving the last inequality we have used Eq. (98,99). Hence to summarize, if the bound in Eq. (100) is violated, then universe will deviate from radiation domination at an intermediate time. On the other hand, if Eq. (100) is satisfied but the bound in Eq. (101) is violated, then radiation domination will be maintained at all times but the RHNs will not immediately decay after being produced from the FOPT. Depending upon the different parameter regions one is interested in, corresponding bounds have to be checked.

### v2

@@ VI Viable UV-complete Particle Theory: Multi-Majoron Model
where $\rm{Br}_{\varphi_{1}\rightarrow N_{3}}$ is the branching ratio of $\varphi_{1}$ into the RHN $N_{3}$ state, and $E_{\max}$ is given in Eq. (43). Note due to the mixing between $\phi_{1}$ and $\phi_{2}$, we have to diagonalize the mass matrix coming from Eq. 76 in order to find $m_{\varphi_{1}}$. Following Ref. Di Bari et al. (2024) we find the corrected mass
@@
Note after $\varphi_{1}$ P.T., $N_{3}$ gets a mass $M_{3}=y_{3}v_{1}/2$. Hence in order for $\varphi_{1}$ to decay into a pair of $N_{3}$ one requires
@@
Hence for our parameter choice with $v_{1}=10^{14}$ GeV, we can neglect the vev correction and simply get an upper bound $y_{3}\lesssim 0.45$. This leads to a corresponding upper bound on $M_{3}$ while calculating different relic density and yields. Note as was discussed earlier, the two lighter RHNs are naturally quasi-degenerate whereas the heaviest RHN exhibits a hierarchical mass scale $M_{3}\gg M_{1,2}$. This near degenerate masses of the RHNs then can increase the CP violation in the framework of resonant leptogenesis Pilaftsis and Underwood (2004). In particular, $\epsilon_{\rm max}\sim\mathcal{O}(1)$ can be taken if $|M_{1}-M_{2}|\sim\Gamma_{1,2}/2$, where $\Gamma_{1,2}$ is the decay rate of the RHNs $N_{1},N_{2}$. This can be achieved by adjusting the corresponding Yukawa couplings $y_{1}$ and $y_{2}$ and hence for the rest of this section we will be taking $\epsilon_{CP}\sim\mathcal{O}(1)$.
@@
Note, in our case $\lambda_{2}/\zeta\lesssim\mathcal{O}(10^{-3})$ hence we can safely assume $\varphi_{2}$ self production from $\varphi_{2}$ bubble collision remains subdominant to $\varphi_{1}$ production. By inspecting Eq. 47 we see the produced $N_{3}$ yield from $\varphi_{2}$ collision is $\propto\ln\left(\frac{2\rm E_{max}}{M_{3}}\right)$ whereas the produced $\varphi_{2}$ yield is $\propto\left[\ln\left(\frac{\rm E_{max}}{m_{\varphi_{1}}}\right)+\frac{8\pi^{2}v_{1}^{2}}{m_{\varphi_{1}}^{2}}\right]$. We note for our benchmark points, $\frac{8\pi^{2}v_{1}^{2}}{m_{\varphi_{1}}^{2}}\gg\ln\left(\frac{\rm E_{max}}{m_{\varphi_{1}}}\right),\ln\left(\frac{2\rm E_{max}}{M_{3}}\right)$, hence the produced $N_{3}$ density always remain subdominant to the produced $\varphi_{1}$ density. Finally, we point out that the mass of the produced $\varphi_{1}$ particles, $m_{\varphi_{1}}\sim\sqrt{2\lambda_{1}}v_{1}\gg v_{2}\sim T_{*}$ for our benchmark and hence the thermal contribution for $\varphi_{1}$ yield is negligible at production.
@@
The corresponding values of the DM relic and BAU for the benchmark points are summarized in Table. 3.Using Eqs. (48) and (49), we find that in the RHN dark matter scenario the observed relic abundance is reproduced for RHN masses in the range $M_{3}\sim 10^{5}$$10^{7}$ TeV. For leptogenesis, the baryon yield $\mathcal{Y}_{B}$ is computed using Eq. (66), assuming $\epsilon_{\rm max}\sim\mathcal{O}(1)$ Pilaftsis and Underwood (2004). For $BP_{Y}$ and $BP_{Z}$ we produce the observed baryon yield given in Eq. 64 for RHN mass $10^{7}-10^{8}$ TeV. On the other hand, for $BP_{X}$, even for the maximum $M_{3}$ allowed by the bound Eq. 88, roughly half of the required asymmetry is produced. This happens due to the relatively low P.T. temperature for $BP_{X}$.
@@
The model also allows for a co-genesis scenario through the inclusion of an additional interaction term of the form $\mathcal{L}\sim y_{\chi}\bar{\chi}\phi_{2}N_{3}$ in Eq. 75. The additional dark sector fermion $\chi$ is charged under $U(1)_{N}\times U(1)_{B-L}$ such that this operator is invariant under the extended multi-Majoron symmetry. The fermion $\chi$ then serves as the DM candidate. The analysis closely parallels that presented in Sec. V. The inclusion of this extra fermionic degree of freedom does not significantly affect the effective thermal potential, since both the thermal and one-loop corrections from fermions are relatively small Salvio (2025). Consequently, the phase transition parameters may be taken to be identical to those computed previously.
@@
In the co-genesis scenario, the dark matter relic abundance is reproduced for $\text{Br}\chi m\chi\sim 5$$50$ GeV. For $\text{Br}_{\chi}\sim 1$, this corresponds to heavy dark matter with $m_{\chi}\sim 10$ GeV; however, in this limit $\text{Br}_{L}\ll 1$ and the resulting baryon asymmetry is negligible. Conversely, for $\text{Br}_{\chi}\sim 0.1$, the observed baryon asymmetry can be obtained simultaneously with a heavy dark matter candidate, $m_{\chi}\sim\mathcal{O}(100)$ GeV.
@@
Figure 15: Parameter scan for a FOPT in the multi-Majoron model in $v_{2}-M$ plane for differnt choices of $\zeta$ and $\lambda_{2}$. Shown are parameter-space points that realize a successful FOPT and produce GW signals with SNR >10>10 at ET and BBO, arising from both bubble collisions and $\varphi_{1}$ production. Regions compatible with successful dark matter production, leptogenesis, and co-genesis are indicated, together with the corresponding values of the RHN mass $M_{3}$. $\kappa_{\varphi_{1}}\sim\lambda\sim\mathcal{O}(1)$, $v_{1}=10^{14}$ GeV and $\lambda_{1}=0.001$ is taken for all the plots.
@@ VII Discussion & Conclusion
  • A non-thermal population of RHNs can be generated from bubble collisions during a FOPT. If the lightest among them is stable then they may constitute the entire DM relic abundance for masses as low as $m_{DM}\gtrsim 10^{6}\ \mathrm{GeV}$ (Fig. 5). Higher transition temperatures correspond to lighter viable DM masses (see Eqs. 48 and 49). The characteristic $f^{1}$ slope from RHN production will be detectable in LISA for $m_{DM}\sim 10^{10}\ \mathrm{GeV}$, SKA for $m_{DM}\sim 10^{14}\ \mathrm{GeV}$ and ET, BBO and future LVK detectors for $m_{DM}\sim 10^{6}\ \mathrm{GeV}$ (see Fig. 6). The novel correlation between the DM relic, mass and the GW amplitude can be understood from Eqs. (34-36,47-49) via $\alpha,\ \beta/H,\ T_{*}$ and $y_{\rm DM}$.
@@ Appendix B Processes leading to Washout of Asymmetries
Here we note down a few washout processes related to our scenario discussed in sec IV. The expressions for all scattering cross sections are approximate and we have dropped the contributions of any thermal masses or subleading logarithmic pieces Cataldi and Shakya (2024) -
@@ Appendix C RHNs and Radiation Domination
Where in deriving the last inequality we have used Eq. (99,100). Hence to summarize, if the bound in Eq. (101) is violated, then universe will deviate from radiation domination at an intermediate time. On the other hand, if Eq. (101) is satisfied but the bound in Eq. (102) is violated, then radiation domination will be maintained at all times but the RHNs will not immediately decay after being produced from the FOPT. Depending upon the different parameter regions one is interested in, corresponding bounds have to be checked.